OpenStudy (anonymous):

is Pb(NO^3)^2(aq)+K^2CrO^4(aq) exothermic reaction or endothermic reaction?

sam (.sam.):

Double replacement reaction gives, \[\begin{array}{c} \text{Pb}\left(\text{N}\text{O}_3\right)_2 \\ \end{array} \text{ + } \begin{array}{c} \text{K}_2\text{Cr}\text{O}_4 \\ \end{array} \text{ $\longrightarrow $ } \begin{array}{c} \text{Pb}\text{Cr}\text{O}_4 \\ \end{array} \text{ + } \begin{array}{cc} 2 & \text{K}\text{N}\text{O}_3 \\ \end{array}\] Using ΔH = energy used in bond breaking reactions − energy released in bond making products To find whether the reaction is exothermic or endothermic. -ΔH gives exothermic reaction, heat is released from the system. +ΔH gives endothermic reaction beacause heat is being added to break the bonds.